An inductor, L, of 2 H and a resistor, R, of 1000Ω are in series across an a.c. voltage of 10 V, frequency 60 Hz. Calculate

the impedance in the circuit
1
Expert's answer
2020-07-20T14:56:41-0400

The impedance in the circuit

"Z=\sqrt{R^2+(2\pi f L)^2}\\
=\sqrt{1000^2+(2\pi \times 60\times 2)^2}=1252\:\Omega"